# MMDesign: Multi-Modality Transfer Learning for Generative Protein Design

### v1

@@ 4.1 Settings
Datasets. The CATH 4.2 [13] dataset is a curated version of CATH [22], in which all available structures with 40% non-redundancy are partitioned by their CATH classification(class, architecture, topology/fold, homologous super-family). The training, validation, and testing splits contain 18204, 608, and 1120 structure-sequence pairs, respectively. The training split is used for training our MMDesign, and the validation and testing splits are for evaluation purposes. CATH specifically refers to CATH 4.2 if not otherwise specified. We also report results on Ts50 & Ts500 [17] to assess the generalization. No canonical training and validation sets exist for Ts50 or Ts500.
@@
Evaluation Metrics. Perplexity and sequence recovery on held-out native sequences are two commonly adopted metrics for protein design tasks. Perplexity measures the inverse likelihood of native sequences in the predicted sequence distribution (lower perplexity for higher likelihood). And sequence recovery measures the accuracy with which the sampled sequence matches the native sequence at each site. To maximize sequence recovery, the predicted sequences are sampled with low temperature $61e-$ 6 from the model.
@@
Implementation Details. During training, we employed the SGD optimizer with a batch size of 5 and a learning rate of 1e-3 to optimize the protein design model. The GVPConv module consists of 4 layers with a dropout of 0.1, and the top-k neighbors are set to 30. The node hidden dimensions of scalars and vectors are 1024 and 256, respectively. For the contextual Transformer module, both the encoder and decoder have 8 layers of 8-head multi-head self-attention blocks, with an embedding dimension of 512 and an attention dropout of 0.1. And 1 NVIDIA A100 80GB GPU is used.
Baselines. Table 1 lists typical and representative works, most of which are open source and easy to re-implement. Works in Group 1 and Group 3 are only trained on a small-scale CATH training set (approximately 18k training pairs). Earlier proposed baselines such as Structured Transformer[13] and GVP-GNN [14] are still competitive and influential with relatively good performances and lightweight advantages. ProteinMPNN [4] by Baker’s group is renowned for improving performance and speed and their impressive biological validation experiments. PiFold[6] follows closely behind with further enhancements in efficiency and effectiveness. ESM-IF[8] by Meta is singled out in Group 2 as a unique open-source representative with strong data augmentation (trained on approximately 12M training pairs from AlphaFoldDB[26]). In addition, our MMDesign’s architectures are similar to ESM-IF, thus ESM-IF is chosen as a baseline.
4.2 Main Results
@@
In line with previous work[13, 14, 8, 6], we start with evaluation in the standard setting of sequence design given all backbone coordinates.
@@
Evaluation on CATH Test Set. Table 1 reports the perplexity (lower is better) and recovery (higher is better) scores on the generic dataset CATH test set (“ALL”). MMDesign achieves the best performance with a perplexity of 3.86 and a recovery of 54.88%, outperforming the state-of-the-art ProteinMPNN, PiFold (Group 1), and ESM-IF (Group 2) by a significant margin.
@@
Evaluation on Single/Short Chains. Following [13], we also evaluate the performance of the models on the subsets of the CATH test set, i.e., “Short” dataset (protein sequence length $\leq$ 100 residues) and “Single-chain” dataset (single-chain proteins recorded in the Protein Data Bank), for evaluation, as shown in Table 1. MMDesign achieves excellent performances on both “Single-chain” and “Short” datasets (although the recovery score of “Single-chain” is slightly lower than GVP-GNN-large on CATH 4.3). In particular, in terms of perplexity, MMDesign (Short: 5.32; Single-chain: 5.43) significantly outperforms PiFold (Short: 6.04; Single-chain: 6.31) and ProteinMPNN (Short: 6.21; Single-chain: 6.68).
@@
Evaluation on Generalization. To fully compare the generalizability of the different models on the out-of-domain datasets, we report the results on the TS50 and TS500 datasets in Table 1. MMDesign still achieves consistent improvements for TS50 and TS500, where the recovery score in Ts50 is close to 60%, and the score in TS500 breaks 61%.
@@
Overall, ESM-IF performs better in generalization on both generic and out-of-domain datasets, while PiFold and ProteinMPNN perform better on short and single-chain datasets. However, MMDesign outperforms all baseline models across a variety of test sets, including generic datasets, out-of-domain datasets, and short and single-chain datasets with balanced performances, showing the strong generalization and robustness of MMDesign. Notably, our results also demonstrate that a pretrained prior knowledge migration-based strategy with small data training can outperform the large data training strategy such as ESM-IF.
@@
Table 2: Ablation study of various pretrained modules based on the protein design framework on CATH validation set and test set (PSM: pretrained structural module; PCM: pretrained contextual module). The best results are bolded.
4.3 Pretrained Module Ablation
To illustrate the significance of the pretrained modules in our MMDesign framework, we conduct an ablation study on the CATH validation set and test set to evaluate the impact of the pretrained structural module (PSM) and the pretrained contextual module (PCM). Overall, both pretrained modules have enhanced effects on the performance, as shown in Table 2.
Comparing #1 vs. #3 and #1 vs. #2, we observe that the pretrained structural features contribute more to the performance. This observation is consistent with our motivation for introducing prior structural knowledge. Moreover, it is likely that the PCM module, as a downstream module of PSM, is directly influenced by the representation capability of PSM.
@@ 4.3 Pretrained Module Ablation
Surprisingly, comparing only #1 vs. #2, we find that performances of MMDesign are improved considerably even without using initialization parameters of the pretrained structural module, i.e., by using only the PCM. Interestingly, this is a completely new observation that has not been extensively explored in previous protein design studies. These findings suggest that the pretrained contextual module has learned rich knowledge of the protein language to improve protein design tasks. Comparing #3 with #4 further confirms this observation.
Figure 4: The designable analysis of different models based on 3D protein structural folding. TM score and GDT score are used to reflect the biologically reasonable performance.
4.4 Generated Sequence Distribution
@@
As shown in Figure 3, we counted both the generated protein residues and raw protein residues according to amino acid residue categories. We generated approximately 181k amino acid residues for each model based on 1,120 protein sequences from the CATH test set. Representative protein design models are chosen for comparison with our MMDesign model. To facilitate comparison, we normalized the distribution using the maximum-minimum normalization process. Compared to the ground-truth distribution (base) derived from raw reference sequences in Figure 3 (green), our MMDesign and PiFold produce distributions that are visually closer to the base. To quantitatively characterize the similarity of the generated sequences to the base distribution, KL-divergence is employed. The results are consistent with the visual observations, where our MMDesign and PiFold have lower KL-divergence values (around 0.03) than other baselines, indicating better similarity to the reference distribution. In contrast, GVP-GNN and Structured Transformer perform poorly, with higher KL-divergence values.
4.5 Designable Analysis
We assume that a biologically reasonable protein design should satisfy the condition that the generated protein sequence can be folded into a designable backbone structure and is physically realizable. Otherwise, the generated protein sequence is considered useless. Building upon previous biological and computational approaches [25, 16, 28], we propose to use ESMFold [18] to fold the generated sequences into their corresponding 3D structures. TM score and GDT score [29] are used to evaluate the structural similarity between the predicted backbone structures against the original native structures.
To our knowledge, we introduce this designable evaluation for the first time to explore the biological feasibility of existing protein design tools. PiFold is excluded from this analysis due to its discarding of coordinate missing nodes, resulting in discontinuous regions that make it difficult to align native structures for a fair comparison with other models. As shown in Figure 4, we first perform a folding procedure on the reference native protein sequences to generate 3D structures that demonstrate a high-precision folding level close to the experimentally validated level (“base” in Figure 4) with a 0.8068 TM score and a 0.7896 GDT score, which is a standard reference here.
We then use selected protein design models (MMDesign, ESM-IF, GVP-GNN, Structured Transformer) to generate sequences on the CATH test set, which are subsequently folded back into 3D structures via ESMFold. Evaluation of the predicted protein backbones reveals that the folding performances of the sequences generated by MMDesign (0.7920 TM score and 0.7755 GDT score) and ESM-IF (0.7916 TM score and 0.7737 GDT score) are significantly better than that of the comparative GVP-GNN and Structured Transformer, approaching the level of “base”, which further demonstrates that the high recovery level of sequences generated by MMDesign is biologically reasonable.
5 Conclusion
@@
The proposed MMDesign represents a significant step forward in protein design by introducing a novel multi-modality transfer learning approach and an explicit consistency constraint. Trained only on the small dataset CATH, MMDesign improves significantly over baselines across various benchmarks. We also systematically present internal and external analysis techniques based on biological reasonable perspectives and data distribution statistics, which greatly improve the interpretability of models and reveal more about the nature and laws of protein design.
In future work, it would be interesting to explore the use of large language models to guide the learning of the alignment between structural and contextual modalities in an end-to-end manner, which could simplify the training pipeline and potentially further improve the performance of MMDesign. Additionally, it would be valuable to evaluate MMDesign on larger and more diverse datasets to further validate its effectiveness and generalizability.

### v2

@@ IV-A Settings
Datasets. The CATH [1, 22] is the main dataset, in which he training, validation, and testing splits contain 18204, 608, and 1120 structure-sequence pairs, respectively. The training split is used for training MMDesign, and the validation and testing splits are for evaluation purposes. CATH specifically refers to CATH 4.2 if not otherwise specified. We also report results on Ts50 & Ts500 [24] to assess the generalization.
@@
Evaluation Metrics. Perplexity and sequence recovery on held-out native sequences are two commonly adopted metrics for protein design tasks. Perplexity measures the inverse likelihood of native sequences in the predicted sequence distribution (lower perplexity for higher likelihood). And sequence recovery measures the accuracy with which the sampled sequence matches the native sequence at each site.
@@
Implementation Details. During training, we employed SGD optimizer with a batch size of 5, a learning rate of 1e-3 to optimize the protein design model, and low temperature 1e-6 for sampling. The GVPConv consists of 4 layers with a dropout of 0.1, and the top-k neighbors are 30. The node hidden dimensions of scalars and vectors are 1024 and 256, respectively. For the contextual Transformer, both the encoder and decoder have 8 layers of 8-head multi-head self-attention blocks, with an embedding dimension of 512 and an attention dropout of 0.1. And 1 NVIDIA A100 80GB GPU is used.
IV-B Main Results
@@
In line with previous work[1, 4, 5, 10], we start with evaluation in the standard setting.
@@
Evaluation on CATH Set. Table I reports the perplexity (lower is better) and recovery (higher is better) scores on the generic dataset CATH test set (“ALL”). MMDesign achieves the best performances with a perplexity of 3.86 and a recovery of 54.88%, outperforming the SOTA ProteinMPNN, PiFold (Group 1), and ESM-IF (Group 2) by a significant margin.
@@
Evaluation on Single/Short Chains. Following [1], we also evaluate the performance of the models on the subsets of the CATH test set, i.e., “Short” dataset (protein sequence length $\leq$ 100 residues) and “Single-chain” dataset (single-chain proteins recorded in the Protein Data Bank), for evaluation, as shown in Table I. MMDesign achieves excellent performances on both “Single-chain” and “Short” datasets (although the recovery score of “Single-chain” is slightly lower than GVP-GNN-large on CATH 4.3). In particular, in terms of perplexity, MMDesign (Short: 5.32; Single-chain: 5.43) significantly outperforms PiFold (Short: 6.04; Single-chain: 6.31) and ProteinMPNN (Short: 6.21; Single-chain: 6.68).
@@
Out-of-domain Generalization. To fully compare the generalizability of the different models on the out-of-domain datasets, we report the results on the Ts50 and Ts500 datasets in Table I. MMDesign still achieves consistent improvements for Ts50 and Ts500, where the recovery score in Ts50 is close to 60%, and the score in Ts500 breaks 61%.
@@
Overall, ESM-IF generalizes better on both generic and out-of-domain datasets, while PiFold and ProteinMPNN perform better on short and single-chain datasets. However, MMDesign outperforms all baselines across a variety of test sets, including generic datasets, out-of-domain datasets, and short and single-chain datasets with balanced performances, showing strong generalization and robustness. Notably, the results also demonstrate that a pretrained prior knowledge migration-based strategy with small data training can outperform the large data training strategy such as ESM-IF.
@@
TABLE II: Ablation study of pretrained modules on CATH validation and test set (PSM: pretrained structural module; PCM: pretrained contextual module). The best results are bolded.
IV-C Performance Analysis
@@
Pretrained Module Ablation. To illustrate the significance of the pretrained modules, we conduct an ablation study to evaluate the impact of the pretrained structural module (PSM) and the pretrained contextual module (PCM). Overall, both have enhanced effects on the performance, as shown in Table II. Comparing #1 vs. #3 and #1 vs. #2, we observe that the pretrained structural features contribute more to the performance. This observation is consistent with our motivation for introducing prior structural knowledge. Moreover, it is likely that the PCM module, as a downstream module of PSM, is directly influenced by the representation capability of PSM. Surprisingly, comparing only #1 vs. #2, performances are improved considerably even without using initialization parameters of the PSM, i.e., by using only the PCM. Interestingly, this is a completely new observation that has not been extensively explored in previous protein design studies. These suggest that the PCM has learned rich knowledge of the protein language to improve protein design tasks. Comparing #3 with #4 further confirms this observation.
@@
Generated Sequence Distribution As shown in Figure 3, we counted both the generated protein residues and raw protein residues according to amino acid residue categories. We generated approximately 181k amino acid residues for each model based on 1,120 protein sequences from the CATH test set. Representative protein design models are chosen for comparison with our MMDesign model. To facilitate comparison, we normalized the distribution using the maximum-minimum normalization process. Compared to the ground-truth distribution (base) derived from raw reference sequences in Figure 3 (green), our MMDesign and PiFold produce distributions that are visually closer to the base. To quantitatively characterize the similarity of the generated sequences to the base distribution, KL-divergence is employed. The results are consistent with the visual observations, where our MMDesign and PiFold have lower KL-divergence values (around 0.03) than other baselines, indicating better similarity to the reference distribution. In contrast, GVP-GNN and Structured Transformer perform poorly with higher KL values.
Figure 4: Recovery score comparison of different models on short, medium and long sequences derived from CATH test set (Length division: Short $\in(0,125]$, Medium $\in(125,188]$, Long $\in(188,500)$. The three divisions are equal in number).
Protein Length Analysis. The comparison on the “Short” dataset in Table I only initially demonstrates the advantage in short sequences. To further clarify the effect of protein sequence length, we divided the CATH test set into equal numbers of short, medium, and long divisions in order of increasing length (“short” here is different from “Short” in Table I). Figure 4 illustrates the comparison with existing advanced baselines on these three divisions. All models perform much worse on short sequences compared to medium and long sequences, consistent with the previous observation in Table I. In terms of short sequences, MMDesign (44.25%) performs slightly worse than PiFold (45.83%) but still outperforms other models greatly. For medium and long sequences, our model has a obvious advantage over others, although ESM-IF does not perform badly either. As expected, the earlier GVP-GNN and Structured Transformer perform worse in all divisions. Overall, our model is relatively better on short sequences and has a significant performance advantage on longer sentences. The comparison provides some insight into the role of the self-attention mechanism in capturing long-range dependencies for longer sequences (PiFold and GVP-GNN do not use self-attention). Although Structured Transformer relies on self-attention for temporal feature representation, it fails to perform well due to the lack of a stronger structural representation module such as GVP.
V Conclusion
@@
The proposed MMDesign represents a significant step forward in protein design by introducing a novel multi-modality transfer learning and an explicit consistency constraint. Trained only on the small dataset, MMDesign improves significantly over baselines across various benchmarks. We also systematically present analysis techniques based on biological reasonable perspectives and data distribution statistics, which greatly improve the interpretability and reveal more about the nature and laws of protein design.
Acknowledgements
This work was supported by National Science and Technology Major Project (No. 2022ZD0115101), National Natural Science Foundation of China Project (No. U21A20427), Project (No. WU2022A009) from the Center of Synthetic Biology and Integrated Bioengineering of Westlake University and Integrated Bioengineering of Westlake University and Project (No. WU2023C019) from the Westlake University Industries of the Future Research Funding.